a cyclist covers half as much as distance as a tanga driver, and the tanga driver half as much again as a man on foot. if man can walk distance of 3 km in one hour how much will cyclist to cover a distance of 27 km
Answer by ankor@dixie-net.com(22740)

a cyclist covers half as much as distance as a tanga driver, and the tanga driver half as much again as a man on foot.
if man can walk distance of 3 km in one hour how much will cyclist to cover a distance of 27 km
:
let c = the cyclist speed (speed and dist are directly related)
"cyclist covers half as much as distance as a tanga driver,"
t = 2c tanga speed (whatever that is)
:
man's speed is 3 km/hr
"the tanga driver half as much again as a man on foot.
t = 1.5(3)
t = 4.5 km/hr
:
Find the cyclist speed
2c = 4.5
c = 4.5/2
c = 2.25 km/hr
"how much time will cyclist require to cover a distance of 27 km"
t = 27/2.25
t = 12 hrs